OpenStudy (anonymous):

Find three consecutive whole numbers such that twice the sum of the two smallest numbers is 35 more than three times the largest number.

OpenStudy (anonymous):

three consecutive numbers are x, x+1, and x+2. So that means: 2(x+x+1)=3(x+2)+35 4x+2=3x+41 x=39 39, 40, 41

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

pleasure. fan me?